### v1

# Evaluating Privacy Perceptions, Experience, and Behavior of Software Development Teams

Abstract
With the increase in the number of privacy regulations, small development teams are forced to make privacy decisions on their own. In this paper, we conduct a mixed-method survey study, including statistical and qualitative analysis, to evaluate the privacy perceptions, practices, and knowledge of members involved in various phases of software development (SDLC). Our survey includes 362 participants from 23 countries, encompassing roles such as product managers, developers, and testers. Our results show diverse definitions of privacy across SDLC roles, emphasizing the need for a holistic privacy approach throughout SDLC. We find that software teams, regardless of their region, are less familiar with privacy concepts (such as anonymization), relying on self-teaching and forums. Most participants are more familiar with GDPR and HIPAA than other regulations, with multi-jurisdictional compliance being their primary concern. Our results advocate the need for role-dependent solutions to address the privacy challenges, and we highlight research directions and educational takeaways to help improve privacy-aware software development.
1 Introduction
The number of privacy violations in the US rose from $\sim$447 in 2012 to $\sim$2100 in 2023 [73], with $\sim$45of companies experiencing harmful breaches [69]. In response, many countries adopted new privacy regulations [46] such as the European General Data Protection Regulation (GDPR) [18]. With these new regulations, developers are under increased scrutiny to implement privacy engineering solutions throughout the Software Development Life Cycle (SDLC) or face financial penalties. Many mobile apps are initially developed by a small team of independent developers with limited privacy expertise or access to legal/policy resources to make privacy decisions [5, 6, 43]. Without mandatory privacy training and a low barrier to entering the market, developers may lack essential knowledge of privacy concepts, leading them to overlook their privacy responsibilities prior to their apps’ deployment.
In recent years, several frameworks and approaches, including Privacy by Design (PbD), are introduced to help developers incorporate privacy rules throughout the SDLC [21, 23, 22, 76, 45, 57, 11, 29, 30, 12, 37, 35]. However, few works examined the implementation of these solutions from the developers’ perspective and their impact on privacy practices. Tahaei et al. [65] suggest that the lack of mainstream adoption of PbD causes confusion for software teams. Recent studies show that developers mainly see privacy as a burden and afterthought and are not familiar with basic privacy concepts [17, 58, 59, 6, 13, 66, 25, 63, 43, 8, 26]. These studies focus on only a limited group of developers and overlook the broader SDLC roles and the unique challenges faced by each specific role. They do not also examine how factors such as legal expertise, privacy regulations, and regional differences influence software teams’ privacy perceptions and practices.
In this paper, we conduct a large mixed-method survey study on Prolific with 189 participants located in the US and 173 participants located in 22 other countries (in total 362), who are involved in various roles in the SDLC – including administrators (e.g., scrum masters, product managers), development and Quality Assurance (QA) teams, and information security/privacy experts. The non-US participants are located in EU+UK (132), South Africa (21), Canada/Mexico (18), and South America (2). Our goal is to identify the current state of privacy comprehension, obstacles, practices, and behaviors in various SDLC roles, and the privacy gaps that have yet to be addressed. Our survey comprises three parts: pre-screening questions (e.g., describing their product/customers), generic questions regarding participants’ demographics (education, role, company size, etc.); and role-specific questions to examine their perceptions, experiences, and behaviors. We combine the participants’ demographics (e.g., location data) with the role-specific responses to help determine:
– RQ1: Are there any differences in privacy perceptions among various roles and demographics (e.g., education, company size, or location)?
– RQ2: Does access to privacy experts (e.g., a Chief Privacy Officer (CPO)) impact privacy perceptions and practices?
– RQ3: How do privacy practices and experiences vary according to SDLC roles, locations, and other demographics?
– RQ4: What is the degree of familiarity of different roles regarding privacy concepts, approaches, tools, and regulations?
To the best of our knowledge, this is the first study to conduct such a holistic evaluation based on the roles in the SDLC.
Our results show that participants have diverse perceptions/definitions of privacy, showcasing the need for a refined approach to privacy in SDLC. Scrum masters, product managers, and information security/privacy experts define privacy more in terms of limited disclosure, while developers and QAs perceive privacy as control over personal information. Our study finds a lack of adoption of most PbD strategies and other privacy techniques, such as Privacy Enhancing Technologies (PETs) and Privacy Impact Assessment (PIA) in SDLC. Most QA members rely on legal/privacy experts to ensure the protection of users’ data, and they lack privacy knowledge and expertise. Members of software teams are generally self-taught regarding privacy concepts, and most are not familiar with regulations that exist in the US, such as the California Consumer Privacy Act (CCPA) [24] and the Children’s Online Privacy Protection Rule (COPPA) [19]. We also find that software teams face challenges in both understanding and adhering to privacy regulations, especially across multiple jurisdictions. These findings highlight the need for more privacy-focused education and training. Comparing regional-specific trends regarding the use of PETs, the creation of PIAs, or the presence of a CPO, we did not observe any differences among our participants regardless of their location. This shows that privacy practices are primarily determined by the culture of the organization, and are not influenced by various regulations across regions [4]. Our results highlight the challenges faced in various SDLC roles and advocate the need for role-dependent solutions to address them. Based on these findings, we outline research directions and educational takeaways to help improve privacy-aware software development.
2 Related Work
Understanding developers’ privacy expertise and concerns has been explored in research through user studies with developers and analysis of developers’ forums [42, 34, 13, 68, 32]. Tahaei et al. [62] conducted 12 semi-structured interviews with developers to examine their strategies to protect users’ privacy. They identified factors such as poor privacy culture, tensions between privacy and other business rules, and lack of standardized privacy tools impact how developers implement privacy. They also emphasize the role of privacy champions to minimize such barriers. Iwaya et al. [34] interviewed 30 practitioners from nine countries, uncovering three key themes: personal privacy mindset, organizational privacy aspects, and privacy engineering practices. They highlight the regulations’ impact on practitioners’ behaviors and corporate cultures, and the need for a privacy-conscious environment. Alomar and Egelman [3] focused on how children’s app developers handle compliance. Through surveys and interviews with 127 participants, they identify that developers often rely on app markets to spot privacy issues and they struggle with third-party tools to maintain compliance. Balebako et al. [6] examined how mobile app developers make privacy and security decisions. Through interviews with 13 developers and a survey of 228 US-based developers, they reveal that smaller companies tend to exhibit fewer positive privacy and security behaviors. Their research emphasizes the need for simplified, cost-effective privacy tools such as privacy checklists, especially for small firms. This study was done in 2014 when regulations such as GDPR and CCPA did not exist. Khandelwal et al. [39] demonstrate that implementing and maintaining privacy labels pose challenges that require resources, tools, and guidelines.
Other works analyze developers’ forums to understand their challenges. The analysis of Stack Overflow (SO) posts shows that developers frequently query regarding PbD, compliance, and confidentiality  [68, 61]. Delile et al. [15] compared privacy questions on SO with responses generated by ChatGPT to identify whether ChatGPT could be used as an alternative tool. Their results indicate in some cases ($\sim$30%), SO is more accurate than ChatGPT responses. Li et al. [41] and Parsons et al. [49] studied posts on several Reddit forums and identified most discussions on personal data usage occur in response to external events such as Android OS changes or privacy laws.
Although our work complements these efforts, it is the first study to assess privacy perceptions, practices, and knowledge of members of software teams involved in various roles in SDLC through a large-scale mixed-method approach. Prior work focused only on developers (i.e., programmers) in the US and a few countries, whereas we study members from various SDLC (including product managers, QA, etc.) roles spanning 23 countries. In our work, we investigate how factors such as organizational aspects (e.g., the presence of a CPO) and participants’ demographics (e.g., role, education, and location) impact privacy perceptions, experience, and behaviors of software teams. We also explore how frequently developers use online forums for privacy-related queries.
3 Study Design
In this paper, we aim to understand how members of software teams in small, medium-sized, and large companies (i.e., with <20, 21–100, and 100+ employees), implement privacy in their software applications and examine their level of privacy comprehension, expertise, practices and behaviors based on various demographics (such as roles, location, education level, etc.). For this purpose, we first conducted a pilot study to evaluate our survey design and then a large-scale study with members of software teams in 23 countries. Our pilot study was completed in January 2023, while our large-scale study was done between February–April 2023.
3.1 Survey Tools
Survey Creation
We utilized Qualtrics for survey creation, a platform supported by our university. Using Qualtrics, we customized our survey to individualize questions based on the participants’ role (Q9) [^1] Survey questions: http://tinyurl.com/2p9n49e4 as defined in Table 1. For example, we asked developers about familiarity with PbD (Q39) and their use of forums such as Reddit (Q32), while information security members were asked about the management of access control, encryption algorithms, and certificates (Q60-Q62).
Table 1: Breakdown of Participants Roles
Role Count
AD: Admin., Product Manager, Scrum Master 70
SD: Software Designer, Architect, Developer 198
QA: Software Tester, Quality Assurance Eng. 40
ISec: Information Security/Privacy Expert 54
Total 362
Survey Platform Selection We conducted the large-scale survey using Qualtrics integration on the Prolific [48, 67] platform since it provides a higher pay rate and allows selecting from a more specific pool of participants with basic programming knowledge, in our case - software teams.
Tahaei et al. and Kaur et al. [67, 38] recommend using Prolific and MTurk for large-scale surveys. Although pre-screening via programming questions is recommended [67, 38, 14, 53], it has limitations: (a) overusing such questions could lead to automatically responding correctly without paying attention to the questions [67]; (b) in studies such as ours where the software teams include a variety of roles (e.g., product manager, QA, etc.) as well as with specific programming skills (e.g., JavaScript developers), having programming knowledge questions may bias the participants’ pools towards more experienced developers in larger companies with traditional programming knowledge, preventing recruiting novice developers and those in other SDLC roles; and (c) AI tools like ChatGPT [10] are widely accessible and can handle code-based questions, thus, these questions are no longer a strong barrier to screen participants. Our analysis of Danilova et al.’s [14] pre-screening questions with GPT-3.5 shows that the tool can answer the questions with 95% accuracy. We discuss how we mitigate these issues below and in Section 4.
Conducting the Survey In our survey, participants were required to be at least 18 years old, fluent in English, and working in industries such as Graphic Design, Information Services and Data Processing, Product Development, Software, Video Games, etc. We paid an average of \$25.17/hr to participants who completed the survey. We filtered participants based on their industry and verified their eligibility on Prolific. Out of the 686 participants who started the survey, 14 did not give their consent, and 295 did not finish the survey; hence, were excluded from our analysis. We then conducted another filtering process, to ensure that the participants work in the software industry and have in fact software development experience. We asked them “In short, tell us about your product and who are your customers.” and manually evaluated their responses. We found that the majority of them are involved in software development activities such as “I make a productivity app for mac & windows to record & share the user’s screen”. We eliminated 15 participants as we could not verify their involvement in SDLC; for example, those with responses as “NA” or “I sell home decor items. My customers are primarily women.” Following these steps, we ended up with a total of 362 participants for our final count.
3.2 Pilot Survey: University Students
Our pilot survey participants were our university’s graduate students over the age of 18 from the disciplines of Computing and Information Science, Electrical and Computer Engineering, and Business, who had experience in software development, IT, or related fields. To maintain their anonymity, we did not collect any personally identifiable information such as their contact, names, or company names.
The goal of the pilot study was to gather initial insights and feedback before the deployment of our main study on Prolific. Upon the IRB approval, we launched the survey using Qualtrics. The survey consisted of 40 questions, including 13 short and 27 multiple-choice questions. We estimated that the survey takes $\sim$30-40 minutes to complete. Every participant was presented with the same set of questions regardless of their role on a software team. We used the responses to improve our large-scale survey (i.e., Subsection 3.3).
We received 45 responses but most were incomplete due to the survey’s length and the diversity of questions. After discussing the study with the participants, we revised and shortened the survey based on participants’ role in the SDLC.
3.3 Software Teams Survey
The main feedback we received from the pilot study was that the survey required too much time to complete ($\sim$27 minutes). To address this limitation and to focus on capturing participants’ perspectives related to their SDLC roles, we separated the survey questions according to the roles. This shortened the survey duration by 12 minutes and enhanced the quality of the responses we received. We first asked all participants the same set of 10 questions that are partly related to demographics and the degree of privacy understanding (e.g., how do they define privacy). We then divided the remainder of the questions into four groups, one for each role defined in Table 1. Our breakdown loosely follows the SDLC phases, but we separated the ISec roles from the SD roles to evaluate the significance of security or privacy knowledge in our survey.
3.4 Survey Questions
The survey includes a mix of demographic, perception, experiential, and behavioral questions to answer our research questions (see Section 1). The complete list of questions (except questions 1-3, which are the required Prolific identification questions and our consent form) is found here. [^2] Survey questions: http://tinyurl.com/2p9n49e4
Demographic Questions These questions collect basic information about the participants such as age, education, their SDLC role, and the company size. e.g.,“What areas/roles of the development team are you currently involved with?”.
Perception Questions These questions aim to understand participants’ perceptions toward privacy; e.g., “How do you define privacy?”.
Experiential Questions These questions ask about the participants’ experience with privacy-related tasks and challenges; e.g., “What was the process for the Privacy Impact Assessment, and who was involved?”.
Behavioral Questions These questions ask about the participants’ behaviors and knowledge related to privacy; e.g., “List any privacy-by-design strategies you have used or know:”.
4 Ethics & Limitations
Ethical Considerations This research adheres to our university’s ethical guidelines and was conducted with our Institutional Review Board (IRB)’s approval. All participants agreed to a thorough consent form that included information about the investigators, the risks, benefits, compensation, and confidentiality. All participants were informed about their voluntary participation, maintaining their right to withdraw at any time. No personally identifiable information was collected; and measures were in place to ensure the anonymity, confidentiality, and security of responses. The contact information of all investigators and the IRB team was also included. No participants contacted the investigators or the IRB about the study or the compensation.
Limitations Like most survey studies, our analysis is based on participant self-report data and is affected by self-report bias, recall bias, and social desirability bias. Participants were informed during consent that the survey pertained to privacy. This may introduce priming and self-selection biases. In addition, there is recruitment bias as the Prolific user base may not fully represent the diverse population of individuals in the SDLC. We used multiple screening questions to ensure that recruited participants in fact have experience in software development activities (Section 3.1). We also asked follow-up and write-in questions to make sure the multiple-choice questions are backed up with written facts. To mitigate the potential for survey responses being generated by AI tools like ChatGPT [10], we minimized open-ended questions in favor of multiple-choice formats and carefully scrutinized the write-in responses to remove those that appeared AI-generated. Short responses with typos and errors suggested that our responses were not AI-generated. Despite our efforts, AI-generated responses could affect the study’s outcomes.
We carefully framed our questions to not prompt biased responses. However, we could not avoid one leading question that asks about the confidence in their companies’ privacy and security measures. We aimed to reduce the bias by providing four options instead of a ‘yes’, ‘no’, with the option to not answer. Additionally, the question follows their own definition of privacy, further helping minimize bias. We employ statistical analyses (like the chi-square test) to ensure the broad applicability of our findings. To control for Type I errors in the presence of multiple hypothesis tests, we report our results after employing Bonferroni correction.
5 Study Analysis Process
Our survey results are organized around our research questions (RQs, see Section 1), with a focus on various areas of privacy within the SDLC and across different roles. Our RQs examine the perceptions held, privacy experience and challenges, and privacy behaviors, while considering the demographic breakdown (see Section 3.4) to provide additional context and to allow for a more nuanced understanding of the data. Our analysis follows a mixed-method approach, encompassing both quantitative and qualitative methodologies.
Qualitative Analysis We evaluate the descriptive and open-ended questions through open coding procedures and iterative processes. For our open-ended questions, e.g., the definition of privacy, the first two authors, independently, classified 50 responses. They then discussed their results and resolved the discrepancies. They continued with the rest of the responses, evaluated the agreements, and resolved the disagreements in another round of discussion. Lastly, a third privacy expert examined the results to ensure their correctness and completeness. For the non-subjective descriptive questions e.g., which PbD strategies they use, one author categorized them and the second author reviewed them, for correctness.
Quantitative Analysis For the questions where our goal is to understand if a correlation exists between the demographics and the privacy-related perception, experience, and knowledge, we conducted statistical analyses. We used Chi-Squared tests to determine whether there is a significant correlation between two categorical variables. For questions where the responses are on a Likert scale, we used the Kruskal-Wallace test. For perception, experience, and behavioral questions, we hypothesize from our RQs that the size of the company, the presence of a CPO or a similar role, the education level, roles, and participants’ location may impact their confidence in privacy/security measures, various privacy practices such as the creation of PIA or privacy policies and their familiarity with PETs, regulations, and usage of forums.
To control Type I errors and avoid false positives, we use Bonferroni correction [54]. Since Bonferroni correction is very conservative and may increase Type II errors, we discuss the results with respect to $alpha=0.05$ as well as the adjusted value (i.e., $\frac{0.05}{24}=0.0021$, for our 24 statistical tests).
6 Findings
6.1 Survey Demographics
In our main study, we received a total of 362 responses (after filtering - see Section 3.1). 189 participants reside in the US and the rest 173 come from 22 other countries (see Section 1). Table 1 shows a breakdown of participants’ roles, with the majority ($\sim$55%) in SD roles. As shown in Appendix A - Table 10, the majority of participants identify as male. Most are below the age of 45 and completed their BSc. (61.05%), with $\sim$61in Computer Science (CS), Information Technology (IT), Data Science (DS), and Electrical & Computer Engineering (ECE) majors. This percentage includes the answers to the “Others, please specify” option. Among those with a Business degree, 61% are in AD roles such as product manager and 28% are in SD roles. Among those in the “Other” degree category, 48% identified as SD, 19.5% as QA, 21.0% as AD, and 11.5% as ISec. Their company sizes of $<$100 employees and 100+ employees are distributed almost equally.
6.2 Perceptions of Privacy
We seek to understand software teams’ privacy comprehension by examining how they define privacy, their confidence in their company’s practices, and if these differ based on roles or organization differences (i.e., RQ1&2).
6.2.1 Definition of Privacy
One of the key questions in our survey is “How do you define privacy?”. The responses were diverse and showed differing perceptions. Some participants defined privacy in terms of data security highlighting the need to protect user data from unauthorized access. For example, one participant explained that “It involves implementing measures to safeguard sensitive information, such as encryption, access controls, and data anonymization". Others described privacy from a user rights perspective. For example, “I define privacy as the ability to control all that is related to my information and to keep it from reaching someone who is unauthorized". Few responses incorporated legal compliance, with one participant defining privacy as: “This involves being compliant with regulations and ensuring all data is protected with a least-privilege access model with ownership of the different part data sources with assigned data stewards".
To effectively categorize the diverse definitions of privacy, we utilized Solove’s taxonomy [56], a comprehensive framework that breaks down privacy into various categories based on the types of harm of a privacy breach. We chose Solove’s taxonomy for two key reasons: (a) it provides a structured and detailed approach to understanding and analyzing definitions of privacy, which is essential with our wide range of definitions and perspectives; (b) it has been widely recognized and used in prominent privacy research [7, 31, 77, 9]. We followed an open coding procedure to map the provided definitions with the taxonomy, as described in Section 5. Multiple classes for each definition were also possible. Figure 1 shows the mapping. (For a breakdown of Solove’s Taxonomy see Appendix B - Table 11; the ‘Blackmail’ category did not apply to any participant’s definitions.)
Figure 1: Privacy Definitions based on Solove’s Taxonomy
Figure 1 shows that the top frequently occurring categories are ‘Disclosure’, ‘Increased Accessibility’, and ‘Insecurity’. This result indicates that most participants either consider the traditional definition of privacy as control over personal information or perceive privacy in terms of security. For ‘Disclosure’, one participant highlights the importance of transparency and clear communication about data collection purposes and user control: “Privacy is the assurance that all data belonging to an individual will be disclosed to others only with that individual’s consent, for uses understood and approved by that individual.” For ‘Increased Accessibility’, a participant, who works with genetic data, underscores the need for controlled access to such information, only granting access if needed: “The users’ ability to define who can access their data and even in that what kind of data can be accessed. As I work in genetic data from patients in my line of work, the clinical information is always controlled access and only researchers working on the particular project can gain access on a need-to-know basis.”
We further examined how privacy perceptions differ across various roles. Almost 50% of participants in AD or ISec roles define privacy as ‘Disclosure’, while QA and SD roles mostly consider privacy as ‘Increased Accessibility’, which is related to access control. ISec roles mentioned ‘Aggregation’ more frequently than other roles, which is an anonymization technique used only in privacy rather than security.
{mdframed}
[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] The variety in our participants’ definitions of privacy shows the complexity of privacy perceptions, and the need for a holistic approach that covers a variety of aspects of privacy throughout the SDLC.
6.2.2 Confidence in Security and Privacy Measures
We asked participants about their confidence in the privacy and security measures implemented in their organization. Table 2 shows the distribution of the participants and their responses. Note that ‘PnS’ stands for ‘prefer not to say’. In all roles, most participants are confident in their company’s security and privacy measures. Interestingly, ISec members are the most confident while the QA members are the most uncertain. This can be either due to QA members considering privacy and security as an afterthought [26], thus ignoring these requirements, or because they encounter more non-compliance instances during testing than any other roles.
Table 2: Distribution of Participants’ Confidence
Role Yes No Unsure PnS
AD 53 (75.7%) 6 (8.6%) 11 (15.7%) 0
SD 149 (75.3%) 8 (4%) 35 (17.7%) 6 (3%)
QA 25 (63%) 2 (5%) 12 (30%) 1 (2%)
ISec 44 (81.5%) 2 (3.7%) 8 (14.8%) 0
Total 271 18 66 7
We analyzed whether there is a correlation between participants’ confidence in security and privacy measures and their demographic factors, such as the company’s size (H1a), participants’ roles (H1b), education level (H1c), and the presence of CPO or a similar position (H1d) (see Appendix C and Table 12 for more details). As shown in Table 12, with Bonferroni adjustment ($\frac{0.05}{24}=0.0021$), we cannot reject the null hypothesis for H1a, H1b, and H1c ($p-value=0.494,0.654$ and $0.570$); thus, we find no correlation between confidence in security and privacy measures and company’s size, participants’ roles, or education levels. However, with a $p-value=0.0007$ for H1d, we can reject the null hypothesis and say there is a correlation between the presence of a CPO (or similar position) and confidence in privacy and security measures. We further evaluate the impact of a CPO below.
We asked the ISec members specific questions regarding their company’s security and privacy measures. When asked “whether their company conducts security audits for third-party software used in their products”, slightly more than half ($\sim$56%) said ’Yes’ while a large number ($\sim$38%) were ‘Unsure’. This is alarming since research shows a large number of third-party software and libraries include security and privacy vulnerabilities [74, 27, 2]. However, when we asked “whether their company securely manages encryption keys and implements encryption algorithms and access control policies”, more than 70% responded ‘Yes’ – which highlights inconsistencies in privacy practices even among experts.
{mdframed}
[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] A CPO is important in fostering employees’ confidence in the privacy and security measures of an organization.
6.2.3 Presence of a Chief Privacy Officer (CPO)
We further evaluate the impact of a CPO or other similar roles on the privacy practices of a company. We focus on the AD and SD roles, who are the majority of our participants (i.e., 268 (74%)). We did not include ISec and QA teams to avoid any response bias due to their active privacy role in the company. We asked them “Do you have a Privacy Officer or similar position in your company?”. Table 13 in Appendix D shows the distribution. Interestingly, only slightly more participants responded ‘Yes’ (42.6%) than ‘No’ (38.4%). $\sim$18were ‘Unsure’, which may indicate the lack of proper communication among employees regarding the company’s privacy practices and the purpose of a CPO. 1.1% responded ‘Other’, which included a legal team or a CTO. Among those that said ‘Unsure’, 23% are in AD roles, and 77% are in SD roles which may indicate CPO members communicate more with the management team (i.e., AD).
Table 3: Distribution of Company Size vs Existence of a CPO
Company Size Yes No Unsure Others
0–20 31.5% 51.5% 15.7% 1.4%
21–100 46.1% 29.2% 21.6% 3.1%
100+ 47.3% 34.9% 17.8% 0%
We investigated whether the larger companies have a CPO. Table 3 shows the distribution of the presence of a CPO based on the company size. Here, we see the presence of a CPO increase with the company size. We also observe that companies of all sizes have a sizable number of ‘Unsure’ responses.
We further asked the participants “When you have a question about compliance with regulations, what do you do?”. The participants could select more than one option. Appendix D - Table 15 shows the distribution of the responses. About half of the respondents (50.1%) mention they ask lawyers or a CPO, while 23.1% look at the best practices and standards (such as NIST guidelines), and 18.5% use developers’ forums (such as Stack Overflow). Among ‘Other’ sources, they mainly mention ‘search Internet’ or ‘ask a colleague’.
We analyzed whether the existence of a CPO (i.e., access to a legal or privacy expert) could impact the creation of PIA (H2a), the familiarity with PETs (H2b), the number of privacy breaches (H2c), or is influenced by the company size (H2d). Appendix D and Table 14 show the list of the hypotheses and the results of the tests. With Bonferroni correction, our results show that the presence of a CPO correlates with the size of a company ($p-value=<0.00001$). This correlation indicates that larger companies are more likely to have a CPO or a legal/privacy expert to help mitigate privacy risks, which is aligned with findings in [6]. However, with the p-value adjustment, we do not find a correlation between the presence of a CPO (or a similar position) and the creation of a PIA ($p-value=0.1005$) and the use of PETs ($p-value=0.008$). This may not be surprising, especially since a majority of SD roles, who are the main users of PETs and are involved in the PIA creation, are unaware of a CPO role. Our analysis also did not reveal a significant correlation between the presence of a CPO and the number of privacy breaches experienced by the organization ($p-value=0.359$). This suggests that the presence of a CPO, while important and necessary, may not be sufficient to help minimize privacy breaches.
{mdframed}
[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] Although a CPO could improve confidence in a company’s privacy measures, it has limited effectiveness in enhancing privacy practices and reducing breaches.
6.3 Experience with Privacy
We ask members of software teams in various roles about their experience with creating privacy policies and/or PIA, as well as practices to ensure the protection of users’ data to better understand their privacy challenges (i.e., RQ2&3).
6.3.1 Creation of a Privacy Impact Assessment
A Privacy Impact Assessment (PIA) is a critical tool for identifying and mitigating privacy risks at any stage of software development. Recently, PIAs and their variations, such as the Data Protection Impact Assessments (DPIAs), have become a requirement in GDPR [18] and CCPA [24]. This tool allows organizations to address privacy and security issues before they become problems. In our survey, we asked participants if a PIA was created at any time throughout development, and if the answer is yes: at what stage it was created, who was involved, and what the process for creation was.
We received 311 responses, where only 43 (14.0%) of them reported that they created a PIA at any point in the SDLC, while a significant proportion (57.2%) reported that they did not (see Appendix E - Figure 5). This indicates a lack of awareness regarding the existence or the need for PIAs (i.e., the PIAs are non-existent or are conducted without their knowledge by the CPO or other teams). We also observed that $\sim$25are unsure about whether a PIA was created, which may highlight a gap in communication within a company about its privacy practices. $\sim$4chose ‘Prefer not to say’.
Among the 43 who created a PIA, 3 of them ($\sim$7%) did not answer the follow-up questions. Our results show that PIAs were created at various stages in the SDLC (see Figure 2), but almost half are at the planning and analysis stages (51.0%). One participant who reported that a PIA was created during the planning stage said, “At the start of development of idea because privacy is more important than all things”. Some participants reported creating a PIA at the start of development, e.g., “We created a Privacy Impact Assessment at the beginning of the software development process. This allowed us to identify potential privacy risks and develop strategies to mitigate them”. Others mentioned during the design phase, or even towards the end of development.
Figure 2: Stages of the SDLC When PIAs are Created.
The sizable number of participants (42%) involved in PIA during the later stages in SDLC may indicate that privacy requirements are not considered early on, and are only included as an afterthought – which is aligned with the findings in [26]. Furthermore, the variation in the timing of the PIA creation shows the need for a more standardized approach to incorporating privacy considerations into software development.
Figure 3 shows the distribution of the roles involved in PIA creation. More than one category was allowed. The responses are also diverse. Some participants reported that they created the PIA themselves or it was a team effort (i.e, SD & QA teams), while others reported that it was done by the CPO or external Legal team, ISec teams, upper management (i.e., CEO or CTO), or even the client (External). This shows that the responsibility for privacy can be distributed across various roles, which again highlights the need for clear communication, collaboration, and defined privacy practice processes.
Figure 3: Distribution of Roles Involved in PIA Creation.
The responses to the question regarding the process for creating the PIA also varied among participants. Some initiate the process by downloading a template and collaborating with internal specialists, while others seek guidance from lawyers, executives, or third-party experts. A common approach involved consulting with professionals, with one participant mentioning that they "outsourced [a] developer that specialises in data privacy and security". Several participants mentioned the involvement of specific roles, such as the Data Chief, IT teams, and privacy protection specialists. The process often involved cross-functional teams. In some cases, senior leadership, such as the CTO, CEO, or owner, played a pivotal role in the assessment.
Lastly, we evaluated the PIA correlation between the company size (H3a) and the participants’ confidence in privacy and security measures (H3b) (see Appendix E). The results show a significant correlation $p-value<0.0001$ for both tests - even after Bonferroni correction.
{mdframed}
[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] Most members of software teams are not familiar with PIA or are unaware of its creation. However, those involved in PIA emphasize the need for its creation in the initial phases of the SDLC in a collaborative process with expertise from various departments and consultants.
6.3.2 Creation of Privacy Policies
Privacy policies describe how, why, and how long an application uses personal information. Regulations [18, 24] and the FTC [70] require companies to provide users with detailed privacy policies. Research shows that these policies may include inconsistencies with apps [76, 55], since they are either created by outside legal experts who may not fully comprehend the apps or by using privacy policy generators [75].
We asked the AD team about their experience and challenges with privacy policies (as other roles are often only indirectly involved). Out of the 70 participants, 3 did not provide any answer. Of the rest, only 11 (17%) have been involved in the creation of a privacy policy, and they used ‘legal experts’ the most (64.0%), followed by ‘templates’ (45.5%), and ‘privacy policy generators’ (36.4%). More than one response could be selected. Two of them mentioned that they either ‘search the Internet’ or ‘ask for team input’, in addition to using privacy policy generators and templates. In 60% (out of 45.5%) of cases that used ‘templates’, and in 50% (out of 36.4%) of cases that used ‘privacy policy generators’, a ‘legal expert’ has also been selected. This result matches with prior research that legal experts in a company are mainly involved in the privacy policy creation, which may lead to inconsistencies between the app and the policy [55]. We also noticed that those who said “Yes” are mostly from companies with less than 100 employees ($\sim$64%) and with a CPO ($\sim$55%).
Finally, we asked the 11 participants who responded ’Yes’, “What challenges did you face when creating your privacy policy?”. We received 10 responses. Six of them describe the challenges regarding compliance with regulations in multiple jurisdictions internationally, and understanding legal jargon, rules, and standards. One specifically had concerns regarding compliance, since they use privacy policy generators: “…differences between different countries and their requirements since we are international.” Five respondents describe their main challenge as ensuring completeness (i.e., covering all personal information), soundness, and language of privacy policies. E.g.: “Whether the wording I chose was going to cover all the bases I needed it to and whether it was clear and easy to understand.” or “Which rules and text to inform users; best guidelines.” One of those five respondents was also concerned about which template to choose. Four others did not find the process challenging since they trusted the legal expert to help.
{mdframed}
[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] Compliance with regulations, and ensuring completeness and correctness are among the most common challenges in creating a privacy policy. Software teams use several tools besides legal experts to help create privacy policies.
6.3.3 Privacy Practices to Protect Users’ Data
We tailored some of the privacy practice questions based on the role, specifically for ISec, SD, and QA teams. We asked ISec members “How do you ensure that data collected from users is used only for intended purposes?”. They discussed various approaches. Some participants (32.08%) emphasized the importance of documentation to ensure transparency and accountability, with one noting “the meticulous documentation of every step in the data usage process”. Encryption emerged as a common theme, with participants mentioning sending encrypted documents and ensuring data is stored securely. A respondent states “I would send documents encrypted and compressed into a zip file, and instruct them to delete the file once the information is accessed.
Limiting access to data is another frequent approach, with 30.2% stressing the importance of restricting data access to only those who need it and maintaining logs to track any access. 16.98% stated the significance of transparency, ensuring they only collect necessary data, obtaining user consent, and regularly monitoring data usage. A few (9.43%) pointed out the importance of adhering to specific regulations, such as the Health Insurance Portability and Accountability Act (HIPAA) [28] and the Family Educational Rights and Privacy Act (FERPA) [44]. A few (16.98%) admitted to not having direct control over data but trusted their organization’s protocols and training to handle data responsibly.
We also asked the same group “How do you manage access to sensitive user data in your organization?”. Role-based access controls, multi-factor authentication, and encryption are common strategies employed to safeguard sensitive information. One respondent shared, “We limit access to systems based on who really needs to access that data.”. Such measures ensure that only authorized personnel can access sensitive data, thereby minimizing potential breaches. Regarding data retention practices, only 47.17% of respondents state that they have been involved in removing user data either after its predetermined lifespan or upon user request.
We asked the SD members “If you encounter a privacy concern at any point in the software development process, what steps would you take?”). More than 95% of them take the concerns very seriously. For example, one participant mentions “run a risk assessment” and another mentions “We take the app offline and start iteratively testing parts of the app to see where the privacy concern is.” About 36% deal with the concern internally to fix it and communicate it with the client and upper management. Another 35% directly escalate it to their supervisors, while 20% seek help from the ISec team or lawyers. A handful contact the client first.
We asked the QA team about third-party systems and their privacy compliance (i.e., “How do you verify that third-party systems used in your products are privacy compliant?”). Similarly, we received diverse responses. Only 56.60% confirmed that their companies conduct security audits of these third-party systems. Some mentioned the significance of conducting vulnerability assessments and penetration testing to ensure third-party systems’ compliance (23%). This is an important approach that ensures external software does not introduce vulnerabilities into the organization’s ecosystem that could lead to a potential breach of privacy if not audited. Some respondents discussed that they rely on reading privacy policies and contracts of third-party systems (28%), while others emphasized the importance of legal agreements and monitoring data transfers (15%). About 22% admitted to not being directly involved in this process, placing trust in their organization’s legal and security teams.
Lastly, regarding the QA teams’ practices for testing for privacy breaches and data leaks, they emphasized the importance of understanding the data they work with and always being vigilant about potential breaches. Regular manual or automated testing is a common theme. $\sim$27of them mentioned the use of penetration testing, both internally and via third-party services. Others stressed the importance of using fake data during testing phases and ensuring that real user data is always encrypted and protected. $\sim$16of respondents admitted to not being directly involved but trusted their organization’s protocols and cybersecurity measures.
{mdframed}
[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] The most common privacy practices among SD, ISec, or QA teams are documentation, auditing, and security techniques (such as access control and encryption). QA teams rely heavily on legal and ISec teams to ensure data protection and are less involved themselves.
6.4 Privacy Awareness and Behaviors
We assess privacy behaviors w.r.t familiarity with regulations, PbD, PETs, and sources of privacy knowledge (i.e., RQ4).
6.4.1 Familiarity with Privacy Regulations
In recent years, several regulations have been introduced that developers need to comply with. Non-compliance with these regulations may lead to financial penalties, sometimes up to 4% of the annual turnover of the company [18]. However, these regulations include legal terminologies that may not be familiar to members of the software teams. To understand the degree of familiarity and awareness, we asked all 362 participants about their familiarity with GDPR, HIPAA, COPPA, CCPA, and the California Privacy Rights Act (CPRA). The answers are on a Likert Scale (see Figure 4).
Figure 4: Familiarity with Different Regulations.
We combined the results from ‘somewhat familiar’, ‘very familiar’, and ‘extremely familiar’ together and found that software teams’ members, regardless of their region and roles, are more familiar with GDPR (77.35%) and HIPAA (63.26%). COPPA, CCPA, and CPRA are all 50% or below. ISec teams are the most familiar with all regulations among all roles, followed by the SD and AD teams. The QA teams are the least familiar with 7.5% familiarity with CCPA and CPRA, and 65.0%, and 57.5% with GDPR and HIPAA.
We asked participants “How did you learn about the previous regulations?”. More than one option could be selected. As shown in Table 4, a majority are self-taught while university education ranks second. Among all roles, the ISec team has the highest percentage of learning about regulations through university education (33.3%), which is more likely through cybersecurity courses. We also asked the participants to describe the other sources they used for learning about privacy regulations. In most cases, they mentioned ‘training at work’ as the source; however, 2 participants mentioned ‘social media’ and ‘YouTube’ as their source. {mdframed}[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] GDPR is the most familiar regulation among all participants due to its comprehensiveness. ISec teams are more likely to learn about regulations through university education; hence, are more familiar with them than other groups. QA teams are the least familiar.
Table 4: Distribution of Participants’ Learning Experience
Role Self Taught Lawyer University Education IAPP Cert. Others
AD 57.1% 5.7% 20.0% 1.4% 15.8%
SD 56.6% 3.5% 21.7% 2.5% 15.7%
QA 85.0% 7.5% 2.5% 0.0% 5.0%
ISec 42.6% 3.7% 33.3% 9.3% 11.1%
Total 57.7% 4.4% 21.0% 3.1% 13.8%
Table 5: Distribution of PbD Strategies Used by Developers
Minimize Hide Separate Abstract Inform Control Enforce Demonstrate
21 22 7 2 17 12 1 4
6.4.2 Familiarity with Privacy by Design (PbD)
Privacy by design (PbD) strategies introduced by Hoepman et al. [30] have gained interest in helping developers to be compliant with regulations. We asked the SD members (i.e., 198 participants) if they are aware of PbD, and if they answered yes, we asked them whether they used them in their work (see Appendix F - Table 17) and to list the ones they used. In general, less than half (46%) are familiar with PbD approaches while $\sim$25are unsure, which indicates the potential knowledge gap and opportunity for educating developers. Out of those who answered ‘yes’ to the awareness of the PbD question, only 57.1% had employed such strategies in their work. Of the remaining, 23.1% did not use them and 16.5% were unsure. This result suggests that even among developers who are familiar with such strategies, not everyone acts on this awareness – which may indicate the lack of usability and readiness of PbD for day-to-day developers’ tasks.
Lastly, we evaluated the responses about the usage of specific PbD strategies (multiple answers were possible). Interestingly, our results are aligned with the findings of Tahaei et al. [64] (see Table 5). Our top categories are ‘hide’ (22), ‘minimize’ (21), ‘inform’ (17), and ‘control’ (12), while ‘enforce’ (1) and ‘abstract’ (2) are rarely discussed. One participant mentions “Mostly minimise. Its the most straightforward.” This response reinforces our result in that ‘minimize’ is one of the easiest strategies to implement. We also received responses regarding Anne Cavoukian’s PbD principles [11] such as ‘privacy by default’ (6 times) and ‘proactive’ (twice). The use of PIA was also mentioned 5 times as a strategy. {mdframed}[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] Our findings show that PbD approaches are not yet commonly used and their lack of adoption underscores the gap in developers’ knowledge regarding PbD and their usability in day-to-day developers’ tasks.
6.4.3 Use of Privacy-Enhancing Technologies (PETs)
Using PETs is another critical component of privacy protection, that allows better protection and maintenance of data privacy against outside threats. We asked the SD team, who are the main users of PETs, if they used any PETs, and if so to list them. Out of the 198 participants, 2 did not respond. 111 of them (56.63%) mentioned that they use some PETs while 36 (18.37%) do not. About 25% are unsure. These results are almost aligned with the degree of familiarity and usage of PbD. There was an increase ($\sim$10%) in PETs familiarity and/or usage in comparison to PbD, which shows that these technologies are more common and tangible for developers, especially those related to encryption and access control. We grouped responses into 8 categories shown in Table 6 (definitions in Appendix G). Encryption and access control, which are primarily security-focused, were the most common, followed by anonymization methods and differential privacy.
Lastly, we investigated the correlations between the PETs’ familiarity and the company size (H4a), confidence in security and privacy measures (H4b), and education level (H4c) (see Appendix H). With the adjusted p-value, we find no correlations ($p-value=0.254$, $0.529$, and $0.704$, respectively).
{mdframed}
[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] PETs are slightly more commonly used than PbD strategies. However, there is still a gap in their familiarity where more than 40% of developers do not use them or are unsure of their usage. The most commonly used PETs are more security-oriented concepts, than privacy.
Table 6: Usage of PETs in Software Development Process
Privacy Enhancing Technology (PET) Percentage
Data Protection and Encryption 70.48%
Access Control 34.29%
Anonymity and Pseudonymity 9.52%
Differential Privacy Approaches 8.57%
Secure Communication 7.62%
Privacy-Enhanced Web Browsing 1.90%
Privacy-Focused Hardware 0.95%
Privacy Governance & Compliance 0.95%
Table 7: Frequency of Usage of the Developers’ Forums
Forums Never Rarely 1-3/M 1-3/W Daily
SO 13.1% 17.1% 26.1% 24.1% 19.6%
GitHub 18.4% 23.9% 23.4% 19.9% 14.4%
Reddit 30.5% 35.0% 20.0% 10.5% 4.0%
Quora 54.5% 27.0% 12.5% 5.5% 0.5%
6.4.4 Developers’ Sources for Privacy Information
As discussed in Section 2, developers sometimes seek privacy-related guidance on forums, such as Reddit or Stack Overflow (SO). We asked the SD teams how often they use various developers’ forums for their privacy-related questions. Table 7 shows the distribution of the responses and their frequencies. $\sim$70and $\sim$58of the respondents use SO and GitHub at least 1-3 times per month, while for Reddit and Quora, this number is about 34.5% and 18.5%. About 57% of the respondents find these forums very or extremely useful, while less than 6% find them not useful at all. In cases where they do not find the answer on these forums, the SD team discusses their questions with the security or privacy experts, asks their teammates, or uses Artificial intelligence tools. In Appendix I, we provide a more detailed analysis regarding the usage of the forums with respect to various factors.
{mdframed}
[backgroundcolor=darkgray!20,linewidth=0pt,innerleftmargin=10pt,innerrightmargin=10pt,skipabove=5pt] Developers often seek privacy-related information from online forums, where more than 50% of participants use either Stack Overflow or GitHub 1-3 times per month and they find these forums useful.
7 Location Analysis
Our large-scale survey included two key regions: the US (189 responses), and non-US (173 responses from 22 countries: EU+UK, South Africa, Canada+Mexico (C+M), and South America), to examine differences in perceptions, experiences, and behaviors. To evaluate the difference in perception, we examine whether participants’ location (US or non-US) correlates with their confidence in privacy and security measures (H6a) and the presence of a CPO (H6b). As shown in Appendix J, both hypotheses do not hold ($p-values$ are 0.1896 and 0.2727). Table 8 shows the presence of a CPO based on the location. The number of ‘Yes’ is almost equal between the US and EU+UK while slightly more US participants mentioned “no CPO” than those in EU+UK. This is not surprising since GDPR, the Data Protection Act of 2018 from the EU, and US HIPAA (Art.164.530) all have a requirement for having a privacy officer or officials in similar role. Interestingly, for South Africa (SA), $\sim$70of the participants said ‘Yes’. The SA Protection of Personal Information Act (POPIA) [1] - Chapter 5 enforces having an information officer.
Table 8: Distribution of Location-based CPO Presence
Locations Yes No Unsure Others
US 43.2% 41.6% 14.4% 0.8%
EU+UK 41.7% 36.1% 20.3% 1.9%
South Africa 69.2% 23.1% 7.7% 0%
C+M 20% 40% 40% 0%
South America 0% 50% 50% 0%
We evaluated whether there is a significant difference between participants’ experience in the two regions regarding the creation of PIA (H6c) and the number of privacy breaches (H6d). With $p-values$ 0.5870 and 0.2016, we find no correlation. We also analyzed the privacy behaviors in the two regions w.r.t familiarity with PbD (H6e) and usage of PETs (H6f). With $p-values$ 0.6246 and 0.2637, we do not find any correlation that suggests that usage of PETs and PbD are equally (un)common in both regions.
Since participants are from regions governed by different privacy laws, we investigated their familiarity with CCPA [24] (H6g) and GDPR [18] (H6h). As expected, we find a significant correlation between the participant’s familiarity with the two regulations (CCPA $p-value<0.0001$, GDPR $p-value<0.0001$). The results for these eight tests do not differ if we compare the US with only the EU and UK.
Due to the global reach of many apps, SDLC teams are responsible for complying with various regulations. We further evaluated the responses to the familiarity with each regulation between the US vs non-US. We combined the responses given for at least somewhat familiarity (i.e., somewhat, very, extremely familiar) and found that participants in the US and South Africa are most familiar with HIPAA while the rest are most familiar with GDPR. Those from US and South Africa are also more familiar with CCPA and CPRA than others. Table 9 shows the distribution. South American participants were not very familiar with any of the regulations.
Table 9: Distribution of Regulations Familiarity
Location GDPR HIPAA COPPA CCPA CPRA
US 71% 85% 54% 49% 45%
EU+UK 89% 37% 38% 11% 9%
S. Africa 73% 77% 73% 46% 46%
C+M 60% 13% 33% 0% 0%
8 Discussion
Summary of Findings With respect to privacy perception, our survey identifies that the majority of the participants define privacy in terms of control over personal information and disclose only when needed, or in terms of security. In other research [32, 62] data protection and security were the most common definitions. Having a CPO or a similar role positively impacts confidence in protecting users’ data. However, we found out that a sizable portion of the participants are unaware of such a role in their company, which may lead to ineffectiveness in utilizing privacy tools or reducing privacy breaches. Lack of proper communication among various roles is a challenge that other research also identified [32, 62]. Our findings also align with Balebako et al.’s 2014 study [6] and Horstmann et al. [32], which observed a correlation between company size and having a CPO. However, we did not observe significant location-based differences in these perceptions. This is interesting but not surprising since GDPR, HIPAA, the UK Data Protection Act, and POPIA all require a CPO or similar roles. Several of our US participants mentioned (in Q27) that they collect Protected Health Information (PHI), which falls under HIPAA; e.g., one participant says “Health related data about people involved with our insurance companies”. The extensive privacy requirements from these regulations likely explain why we observed no significant geographical differences in terms of participants’ confidence, familiarity with PbD, and the usage of PETS.
In terms of privacy experience, most participants rely on legal experts to help create privacy policies; unlike [6] where creating a privacy policy was not the priority. Our study also shows that participants are primarily concerned about multi-jurisdictional compliance. Most of them are not involved in creating a PIA. The majority of those involved believe a PIA should be created during the planning or analysis phases, but some still create a PIA during development, testing, and deployment; this is almost similar to findings in [26, 32]. Our participants emphasized the importance of detailed documentation regarding the data lifecycle, as well as using encryption and access control tools to protect the confidentiality and integrity of data. Interestingly, the QA teams rely more than others on security, privacy, and legal experts to implement and enforce privacy and security rules. Other studies did not examine the privacy practices of QA roles, separately.
Regarding privacy behavior, we identified that less than half of the participants are aware of PbD and an even smaller number use them. Similar to[64], ‘hide’, ‘minimize’, ‘inform’, and ‘control’ are more commonly used. The usage of PETs is slightly more prevalent than PbD, but the focus is more on security practices, such as encryption and access control; similar to other research that found security concepts are more tangible [26, 6, 62, 32]. Anonymization techniques are not used frequently enough. We also find that although $\sim 52\%$ of our participants are from the US, most are more familiar with GDPR than US-based regulations such as COPPA and CCPA. ISec experts are among the most knowledgeable about various regulations, while QA teams are the least familiar. Other work focus mainly on GDPR and CCPA and do not explore details regarding participants’ familiarity [26, 6, 62, 32]. Most participants tend to seek answers to their privacy questions from developers’ forums in addition to legal/policy experts; unlike [6] where they used ‘friends’ or ‘social media’.
Research Directions Insights from the related work and our survey results highlight the need for approaches to operationalize PbD strategies and incorporate them into design and development. PbD patterns [71, 72] provide detailed information about their usage and high-level solutions but still lack implementation. Approaches that detect privacy behaviors in code [37, 35] and further link them to patterns or leverage automated code generation techniques to generate code from privacy patterns are yet to be explored.
Our survey highlights software teams’ challenges while creating PIAs and privacy policies. Research directions that focus on automated approaches to detect the information types, privacy practices, and purposes pre- [33] and post-development [37, 35], or to generate privacy statements directly from code [36] could aid software teams in creating more accurate PIA and privacy notices – which alleviate the challenges regarding consistency and compliance.
Developers seek answers to their privacy-related questions from developers’ forums, though increasingly use tools such as ChatGPT [10, 47]. However, these tools may not always provide accurate responses [15]. Developing methods to help translate developers’ privacy-related questions into accurate privacy code snippets requires further attention [20].
Our survey indicates that software teams face challenges in understanding and adhering to privacy regulations; thus, there is a need for approaches to help better understand such regulations, and establish and maintain compliance. However, most research focuses on detailed requirements analysis, not suitable for agile app development. Future studies could focus not only on automated extraction of legal/privacy requirements but also on generating (privacy-related) user stories to be used in agile development. Research directions on automated approaches to monitor compliance and nudge developers towards compliant approaches are also worth addressing.
Educational Takeaway Similar to other work [6, 32, 62], our work shows the need for a more focused educational approach toward privacy in the SDLC. While currently, many courses emphasize security, it is important to tailor specific courses that include advanced privacy topics such as: regulations; the importance of PIA and other artifacts; challenges in notice, choices, and privacy policy creation; and approaches such as PbD, differential privacy, and federated learning. This distinction between privacy from security is crucial since privacy encompasses a broad spectrum of concerns, including data handling, user consent, and transparency.
Software teams should be equipped with educational modules and tool supports that foster and support life-long learning of dynamic privacy concepts. Nudging developers towards more privacy-preserving solutions through online support and tools is important. Balebako et al. [5] suggest that with the right guidance, developers can be encouraged to prioritize privacy in their design and development processes.
9 Conclusion
In this paper, we examined privacy perceptions, practices, and awareness of SDLC team members during software development. Our findings suggest a need for standardized privacy practices, educational awareness and implementation of PbD, and a privacy expert to promote privacy awareness and compliance. We identified gaps in privacy practices among software teams. Finally, we provide research and educational directions to reduce the challenges in implementing these practices.
In the future, we will extend our research to conduct a comparative analysis within the US states. We will also evaluate whether developers over-claim their expertise in a new study. We will look into how privacy is taught at educational institutes, both in computer science and at Law schools.
Appendix A Participants’ Demographic Information
Table 10 below shows the various demographics of our participants.
Table 10: Demographic Information about the Participants
Gender Female (25.48%) Male (73.41%) Non-Binary (0.55%) Other (0.55%) PnS (0%)
Age 18-25 (19.89%) 26-35 (45.86%) 36-45 (20.99%) 46-55 (8.84%) $>$55 (3.87%)
Education High school (10.22%) BSc. (61.05%) MSc. (22.10%) PhD (1.66%) Other (3.87%)
Degree CS/ECE/DS (34.8%) IT (26.24%) Business (11.05%) Other (24.04%) PnS (3.87%)
Company Size 100+ emp. (50.00%) 50-100 (13.54%) 21-50 (12.43%) 11-20 (7.46%) 0-10 (16.57%)
Appendix B Details of Solove’s Taxonomy
Solove’s Taxonomy and the mapping of subcategories.
Table 11: Solove’s Categories and Subcategories
Main Category Solove’s Subcategories
Information Collection Surveillance, Interrogation
Information Processing Aggregation, Identification, Insecurity, Secondary Use, Exclusion
Information Dissemination Breach of Confidentiality, Disclosure, Exposure, Increased Accessibility, Blackmail, Appropriation, Distortion
Invasion Intrusion, Decisional Interference
Appendix C Confidence in Security & Privacy Measures
The hypotheses list for the correlation between confidence in security and privacy measures and various factors are:
– H1a: The size of the company correlates with confidence in privacy and security measures.
– H1b: The participants’ role at the company correlates to confidence in privacy and security measures.
– H1c: The education level correlates to confidence in privacy and security measures.
– H1d: The presence of a CPO or similar position correlates to confidence in privacy and security measures.
The p-value results of the Chi-Square tests are as follows:
Table 12: P-Value for Hypothesis H1a to H1d
H1a H1b H1c H1d
P-Value 0.494 0.654 0.570 0.0007
Appendix D Presence of a CPO or a Similar Role
The participant’s knowledge about the presence of a CPO in their company is as follows:
Table 13: Distribution of Knowledge about a CPO
Yes No Unsure Others
42.6% 38.4% 17.9% 1.1%
The list of the hypotheses for the correlation between the presence of a CPO or a similar role and the creation of a PIA, familiarity with PETs, number of privacy breaches, and the size of a company are as follows:
– H2a: The creation of a PIA correlates to the presence of a CPO or similar position at the company.
– H2b: Familiarity with PETs correlates to the presence of a CPO or similar position at the company.
– H2c: The higher number of privacy breaches correlates to the presence of a CPO or similar position at the company.
– H2d: The size of a company correlates to the presence of a CPO or similar position at the company.
The p-value results of the Chi-Square tests are as follows:
Table 14: P-Value for Hypothesis H2a to H2d
H2a H2b H2c H2d
P-Value 0.1005 0.008 0.359 $<0.00001$
The distribution of how participants address their compliance questions:
Table 15: Distribution of Sources for Compliance Questions
Lawyer CPO Best Practices Forums Others
24.2% 25.9% 23.1% 18.5% 8.3%
Appendix E The Creation of a PIA
The list of the hypotheses for the correlation between the creation of a PIA and the size of a company and confidence in privacy and security measures are as follows:
– H3a: The size of the company correlates to the PIA creation.
– H3b: The participants’ confidence in an organization’s privacy and security measures correlates to the PIA creation.
The p-value results of the Chi-Square tests are as follows:
Table 16: P-Value for Hypothesis H3a to H3b
H2a H2b
P-Value $<0.00001$ $<0.00001$
The distribution of responses to the creation of a PIA in their company is shown below.
Figure 5: Distribution of Responses to the Creation of a PIA.
Appendix F Privacy by Design Approaches
The distribution of participants who are familiar with PbD:
Table 17: Distribution of Familiarity with PbD Strategies
Role Yes No Unsure PnS
SD 91 (46%) 54 (27.3%) 49 (24.7%) 4 (2%)
Appendix G The Most Commonly Used PETs
Data Protection and Encryption: A widely adopted technique, encryption converts data into a code to prevent unauthorized access. It is used for both data at rest (stored data) and data in transit. Examples of encryption techniques mentioned include AES-256 [51] and SSL/TLS [50].
Access Control: This involves implementing policies that restrict access to sensitive data to authorized individuals only. Role-based access control, where users are granted access to specific features based on their roles in the organization, is a common practice [52].
Anonymity and Pseudonymity: These techniques involve removing personally identifiable information (PII) to prevent individual users from being identified. Pseudonymity involves replacing identifiers with pseudonyms [60].
Differential Privacy Approaches: This technique involves adding noise to the data to protect individual user information while still providing useful insights. It is particularly useful in data analysis and machine learning applications [16].
Secure Communication: This involves encrypting all communications within the software using standard protocols like HTTPS and SSL/TLS [50].
Privacy-Enhanced Web Browsing: This method uses technologies like tracker blocking, web request encryption, and privacy-centric search engines to enhance online privacy. It minimizes the personal data third parties can gather during web browsing.
Privacy-Focused Hardware: This refers to hardware built with privacy at its core, featuring physical switches to disable microphones and cameras, and often running on open-source software. It aims to give users control and protect their data from unauthorized access.
Privacy Governance and Compliance Technologies: These systems help organizations comply with data protection laws like GDPR and HIPAA by assessing, managing, and reporting data privacy practices.
Privacy-Focused Operating Systems and Tools: They are built to prioritize user privacy and reduce data exposure. They usually feature pre-installed privacy tools and default encryption. They might include messaging apps with end-to-end encryption or services that avoid tracking user data.
Appendix H Detailed Analysis of PETs’ Familiarity
The list of the hypotheses for the correlation between the usage of PETs and the size of the company, participants’ confidence, and the presence of the CPO is as follows:
– H4a: The size of the company correlates to the use of PETs.
– H4b: The participant’s confidence in an organization’s privacy and security measures correlates to the use of PETs.
– H4c: The participant’s education level correlates to the use of PETs.
Table 18 shows the results of the hypotheses analysis.
Table 18: P-Value for Hypothesis H4a to H4c
H4a H4b H4c
P-Value 0.254 0.704 0.529
Appendix I Factors Influencing Usage of Forums
To further evaluate the impact of the size of the company, familiarity with PETs, and the presence of a CPO on the usage of developer forums, we employed the Kruskal-Wallis test which is a non-parametric test that is used to compare two or more independent samples for statistically significant differences between groups [40]. Below is the list of hypotheses for the frequency of the usage of the developers’ forums:
– H5a: The size of the company correlates to the use of developer forums to ask privacy-related questions.
– H5b: The presence of a Chief Privacy Officer or similar position at a participant’s organization correlates to the use of developer forums to ask privacy-related questions.
– H5c: Familiarity with PETs correlates to the use of developer forums to ask privacy-related questions.
As shown in Table 19 when comparing forum usage with the size of the company, a statistically significant difference was found between the groups ($H-Value=4.03,p-value=0.04$). However, no significant difference was noted when comparing forum usage with the presence of a Chief Privacy Officer (CPO) ($H-value=9.83,p-value=0.08$). No significant differences were detected when comparing forum usage with the usage of Privacy Enhancing Technologies (PETs) ($H-value=3.92,p-value=0.56$). These findings suggest that the size of the company and the presence of a CPO are more likely to influence the frequency with which developers consult forums for privacy-related inquiries, while the use of PETs appears to have less impact on this behavior.
Table 19: P-Value and H Value for Hypothesis H5a to H5c
H5a H5b H5c
P-Value 0.04 0.17 0.08
H Value 4.03 7.80 9.83
Appendix J Details for the Location Analysis
Below is the list of hypotheses for location analysis. and
H6a: The participants’ confidence in their organization’s privacy and security measures correlates to their region of origin.
– H6b: The presence of a CPO or similar position at a participant’s organization correlates to their region of origin.
– H6c: The participants’ creation of a PIA correlates to their region of origin.
– H6d: The participants’ organization being a victim of a breach of privacy correlates to their region of origin.
– H6e: The participants’ familiarity with PbD strategies correlates to their region of origin.
– H6f: The participants’ use of PETs correlates to their region of origin.
– H6g: The participants’ familiarity with the CCPA correlates to their region of origin.
– H6h: The participants’ familiarity with the GDPR correlates to their region of origin.